### v1

# CAD-VAE: Leveraging Correlation-Aware Latents for Comprehensive Fair Disentanglement

Abstract
While deep generative models have significantly advanced representation learning, they may inherit or amplify biases and fairness issues by encoding sensitive attributes alongside predictive features. Enforcing strict independence in disentanglement is often unrealistic when target and sensitive factors are naturally correlated. To address this challenge, we propose CAD-VAE (Correlation-Aware Disentangled VAE), which introduces a correlated latent code to capture the shared information between target and sensitive attributes. Given this correlated latent, our method effectively separates overlapping factors without extra domain knowledge by directly minimizing the conditional mutual information between target and sensitive codes. A relevance-driven optimization strategy refines the correlated code by efficiently capturing essential correlated features and eliminating redundancy. Extensive experiments on benchmark datasets demonstrate that CAD-VAE produces fairer representations, realistic counterfactuals, and improved fairness-aware image editing. Code
Figure 1: Generated Fair Fine-Grained Image Editing samples from CAD-VAE. The blue-framed section interpolates $z_{Y}$ and $z_{S}$ while fixing $z_{R}$ and $z_{X}$. The red-framed section interpolates $z_{Y}$ and $z_{R}$, replacing $z_{S}$ with the reference while keeping $z_{X}$ constant. $z_{R}$ enables fine control over sensitive relevant features like makeup or mustache. See details in Section 4.3
1 Introduction
Deep generative models have achieved remarkable success in capturing complex data distributions for applications ranging from image synthesis [1] to video generation [34]. In particular, variational autoencoders (VAEs)[13, 21, 31, 33] have provided a principled approach to representation learning, where data are encoded into compact latent variables that effectively capture meaningful factors of variation. However, while these latent representations have enabled impressive performance in numerous tasks, concerns about fairness have emerged, as models can inadvertently learn and amplify biases present in training data[15].
Such fairness issues arise when the target label and sensitive label become entangled due to societal or dataset biases [24, 26]. To address these problems, existing methods commonly fall into two categories. Invariant learning techniques aim to remove sensitive attributes from the learned representation, often via adversarial training or additional regularization [24, 40]. By contrast, disentanglement approaches encourage the model to partition its latent space into separate codes for target and sensitive information, seeking statistical independence among them [6, 28]. Although these solutions have made progress, they typically assume minimal correlation between target and sensitive factors or enforce strict separation via mutual information penalties [3]. However, multiple works [15, 36] have demonstrated that achieving fully fair disentanglement is fundamentally impossible under realistic conditions. First, many datasets contain unwanted correlations between the target label and sensitive attributes due to societal bias, making it infeasible to preserve all predictive cues while completely discarding sensitive information [8]. Second, certain features inherently influence both target and sensitive attributes, so perfectly partitioning features into disjoint latent spaces is unachievable without compromising prediction accuracy [22]. In these circumstances, any attempt at full disentanglement faces an inevitable trade-off between fairness and utility.
A natural way to handle this correlation is to explicitly model how target and sensitive attributes overlap. For instance, some methods rely on causal graphs to separate task-relevant features and capture their relationships with sensitive variables [20, 41, 51, 14]. However, constructing such graphs requires extensive domain knowledge, which is often challenging to acquire in real-world scenarios.
Motivated by the limitations of existing methods, we introduce an additional latent code specifically designed to capture the shared information between target and sensitive attributes. Our approach directly minimizes the conditional mutual information between target and sensitive property predictions, conditioned on this correlated latent code. This design guarantees that the target and sensitive latent codes remain independent once the overlapping factors are absorbed. Moreover, an explicit relevance learning strategy ensures that the correlated latent code captures only the essential shared information without extra domain knowledge. In summary, our contributions are as follows:
  1. 1.
    We propose a novel correlation-aware representation learning framework that directly minimizes the conditional mutual information between target and sensitive property, conditioned on the correlated latent code, effectively addressing the conflict between predictive objectives and disentanglement.
  2. 2.
    We introduce an explicit relevance-driven optimization strategy that precisely regulates the correlated latent code, ensuring it captures only the essential shared information without extra domain knowledge.
  3. 3.
    We validate our approach through comprehensive experiments on multiple benchmark datasets, demonstrating its superiority in achieving correlation-aware disentanglement, enhancing fair prediction performance, and improving both counterfactual generation and fairness-aware image editing, as well as its broad applicability in the context of Vision-Language Models (VLM).
2 Related Work
2.1 Fair Disentanglement Learning
Instead of directly learning sensitive-information-free representations [28, 30, 46], Fair Disentanglement Learning focuses on disentangling these representations into two sets of latent variables, each dedicated to capturing target or sensitive information separately [43]. One of the pioneering works, $\beta$-VAE [13], introduced a framework for semantic feature disentanglement by enforcing a stronger prior constraint on the variational distribution. Extending this idea, FactorVAE [19] proposed minimizing the total correlation (TC) among latent variables to achieve a more refined decomposition of latent space. Specifically, FactorVAE introduces a discriminator-based approach to approximate the TC loss as:
$L_{TC}=KL(q(z)||\prod_{j}q(z_{j}))\approx\mathbb{E}{q_{\phi}(z)}\left[\log\frac{D(z)}{1-D(z)}\right],$ (1)
where the discriminator $D$ aims to distinguish whether the latent code $z$ is sampled from the aggregate posterior or from a combination of in-batch permutations of marginal distributions across each latent dimension $j$. Meanwhile, the encoder is trained to fool the discriminator, promoting independence among latent dimensions. Further advancements on FactorVAE, FairFactorVAE [28] incorporate strategies to obfuscate sensitive information within the latent space, enhancing privacy preservation.
Another notable approach, FFVAE [6], extends the use of total correlation [3] specifically to disentangle the sensitive code with adaptability to the dimension of sensitive information. GVAE [7] employs adversarial training to minimize the leakage of unwanted information in each latent code, offering an alternative method to enforce independence between target and sensitive representations. Meanwhile, ODVAE [42] introduces orthogonal priors for target and sensitive codes, ensuring that these codes maintain minimal overlap. FairDisCo [27] approaches the independence challenge through distance covariance minimization, offering a non-adversarial alternative to traditional methods.
2.2 Correlation-Aware Learning
While disentanglement methods advance different information separations, achieving perfectly independent latent codes remains challenging. The need to recover both sensitive and target information often conflicts with maintaining complete independence, as these types of information are frequently correlated in practice [32, 15]. For instance, in facial attribute recognition on the CelebA dataset [29], ’mustache’ (sensitive relevant) correlates with both gender (sensitive) and attractiveness (target), highlighting the inherent dependency between sensitive and target codes.
To address the challenge of correlated information, Correlation-Aware Learning [20, 41, 51, 14] employs a causal graph approach to separate latent variables based on their correlation with sensitive information. However, constructing such graphs requires domain knowledge and may not always be feasible. Additionally, not accounting for causal relationships between variables can hinder the independence of learned latent codes, leading to practical and computational challenges [15].
To address the inherent correlation between target and sensitive information while reducing dependence on domain knowledge, FADES [15] proposes a novel approach that minimizes conditional mutual information by grouping samples based on different attributes to learn shared information in the sensitive relevant code. Although innovative, this indirect method may not effectively eliminate unwanted information leakage and lacks explicit guidance to ensure the sensitive relevant code captures an appropriate and balanced amount of information. These limitations motivate our approach, which directly minimizes conditional mutual information to achieve complete disentanglement and provides explicit constraints to ensure accurate relevance capture.
2.3 Counterfactual Fairness
Counterfactual fairness (CF) ensures fairness by evaluating whether predictions remain consistent when sensitive attributes are altered while keeping other features unchanged, aiming to mitigate bias through comparisons between factual and counterfactual outcomes [23]. The causal inference technique is widely adopted to generate these counterfactual instances [50, 16, 51, 4, 45], enabling a comparative analysis between actual and hypothetical scenarios.
Graph-based models with predefined intervention and individual variables [20, 25] use causal effect analysis to capture feature interactions for realistic counterfactuals. However, their effectiveness relies on precise domain knowledge; without it, generated counterfactuals may become unrealistic or nonsensical, such as depicting a girl with a mustache.
While CF demonstrates strong theoretical foundations, its dependence on precise domain knowledge complicates model design and may limit practical utility. To address this, our approach introduces a correlated latent code with an explicit relevance learning strategy, allowing the model to learn attribute relationships autonomously and enhance counterfactual fairness without external knowledge.
3 Method
3.1 CAD-VAE
Let $\mathcal{D}=\{(x_{i},y_{i},s_{i})\}_{i=1}^{N}$ denote a dataset consisting of triplets, where $x_{i}$ denotes an input sample (e.g., an image), $y_{i}$ is the label of $x_{i}$ corresponding to target property $Y$, and $s_{i}$ is the label corresponding to the sensitive property $S$. The value range of $Y$ and $S$ is $\mathcal{Y}$ and $\mathcal{S}$, respectively, i.e., $y\in\mathcal{Y}$ and $s\in\mathcal{S}$. The goal is to learn a latent representation that factorizes the information pertinent to $Y$, the information pertinent to $S$, the shared information between $Y$ and $S$, and the background or irrelevant factors. We introduce the following latent representations to capture different information:
  • $z_{X}$: captures task-irrelevant information.
  • $z_{Y}$: encodes the information strongly correlated with $Y$.
  • $z_{S}$: encodes the information strongly correlated with $S$.
  • $z_{R}$: represents the shared information between $Y$ and $S$.
Hence, for a single observation, the corresponding latent variable set is $z:=(z_{X},z_{Y},z_{S},z_{R})$. We first present the problem definition, model components, and architecture in Section 3.1, which serve as crucial foundations for the subsequent sections. To achieve correlation-aware disentanglement learning, we propose directly minimizing the conditional mutual information between $z_{Y}$ and $z_{S}$ with respect to their corresponding opposite attributes $S$ and $Y$, conditioned on $z_{R}$. This approach is complemented by an explicit relevance learning strategy that constrains $z_{R}$ to effectively capture shared information between $Y$ and $S$ while avoiding redundant information. Detailed explanations of these strategies are provided in Section 3.3 and Section 3.4, respectively.
To learn such latent code, we employ a Variational Autoencoder (VAE) [13] with an encoder $q_{\phi}(z\mid x)$ and a decoder $p_{\theta}(x\mid z)$ as our backbone. The prior distribution over the latent variables $p_{\theta}(z)$ is typically chosen to be factorized as a standard Gaussian: $p_{\theta}(z)=\mathcal{N}(z\mid 0,I)$. To train the VAE, we minimize the negative evidence lower bound (negative ELBO), defined as
$\min_{\theta,\phi}\Bigl{[}\mathcal{L}_{\text{VAE}}(\theta,\phi)\Bigr{]}\;=\;\mathbb{E}_{q_{\phi}(z\mid x)}\bigl{[}-\log p_{\theta}(x\mid z)\bigr{]}$ (2)
$\;+\;\mathrm{KL}\Bigl{(}q_{\phi}(z\mid x)\,\big{\|}\,p_{\theta}(z)\Bigr{)},$
where $\theta$ represents the parameters of the decoder and $\phi$ represents the parameters of the encoder.
In addition, we introduce four classifiers to enforce different constraints, including:
  • Eliminating information leakage (see in Section 3.3)
  • Encouraging $z_{R}$ encapsulate only the correlated information $Y\cap S$ (see in Section 3.4)
  • Enforcing $z_{Y}$ and $z_{S}$ to capture sufficient information ensures that attributes $Y$ and $S$ can be recovered correspondingly in alignment with $z_{R}$.
Here, we first present the training method and loss function of each classifiers.
  • $f_{y}(z_{Y},z_{R})$ is a classifier that predicts $\hat{y}$ from $(z_{Y},z_{R})$;
  • $f_{s}(z_{S},z_{R})$ predicts $\hat{s}$ from $(z_{S},z_{R})$;
  • $f_{y\_op}(z_{S})$ is an opponent classifier that attempts to predict $\hat{y}$ from $z_{S}$;
  • $f_{s\_op}(z_{Y})$ attempts to predict $\hat{s}$ from $z_{Y}$.
Let $\omega_{y}$, $\omega_{s}$, $\omega_{y\_op}$, and $\omega_{s\_op}$ denote the parameters of these four classifiers, respectively.
We define
$\min_{\phi,\omega_{y}}\Bigl{[}\mathcal{L}_{y}(\omega_{y},\phi)\Bigr{]}\!=\!\mathbb{E}_{(x,y)\sim\mathcal{D}}\Bigl{[}-\log f_{y}\bigl{(}\hat{y}\,\big{|}\,z_{Y},z_{R}\bigr{)}\Bigr{]},$ (3)
where $z_{Y}$ and $z_{R}$ are sampled from the encoder $q_{\phi}(z\mid x)$: $(z_{Y},z_{R})\sim q_{\phi}(z\mid x)$. The parameters $\omega_{y}$ and the encoder parameters $\phi$ are jointly updated to reduce the cross-entropy in (3), ensuring that $(z_{Y},z_{R})$ carry sufficient information about $Y$. Similarly, the classifier $f_{s}(z_{S},z_{R})$ predicts $s$:
$\min_{\phi,\omega_{s}}\Bigl{[}\mathcal{L}_{s}(\omega_{s},\phi)\Bigr{]}\!=\!\mathbb{E}_{(x,s)\sim\mathcal{D}}\Bigl{[}-\log f_{s}\bigl{(}\hat{s}\,\big{|}\,z_{S},z_{R}\bigr{)}\Bigr{]}.$ (4)
To measure the information leakage, which is illustrated in Section 3.3,we introduce:
$\min_{\omega_{y\_op}}\Bigl{[}\mathcal{L}_{y\_op}(\omega_{y\_op};\phi)\Bigr{]}\!=\!\mathbb{E}_{(x,y)\sim\mathcal{D}}\Bigl{[}-\log f_{y\_op}\bigl{(}\hat{y}\,\big{|}\,z_{S}\bigr{)}\Bigr{]},$ (5)
where $z_{S}\sim q_{\phi}(z\mid x)$ is produced by the frozen encoder i.e $\phi$ is not updated during the minimization of (5); this network is trained to detect any $Y$-relevant information that may unintentionally exist in $z_{S}$. Analogously, the classifier $f_{s\_op}(z_{Y})$ aims to predict $s$ given $z_{Y}$:
$\min_{\omega_{s\_op}}\Bigl{[}\mathcal{L}_{s\_op}(\omega_{s\_op};\phi)\Bigr{]}\!=\!\mathbb{E}_{(x,s)\sim\mathcal{D}}\Bigl{[}-\log f_{s\_op}\bigl{(}\hat{s}\,\big{|}\,z_{Y}\bigr{)}\Bigr{]}.$ (6)
Likewise, $\phi$ is fixed, and only $\omega_{s\_op}$ is updated when minimizing (6).
3.2 Conditional Independence
As discussed in the Related Works, correlated information between target attribute $Y$ and sensitive attribute $S$ is pervasive in disentanglement learning tasks. To address this challenge, we introduce an additional latent code $z_{R}$ to capture this correlated information, as described in Section 3.1.
The latent code $z_{R}$ specifically directs the overlapping information between $Y$ and $S$, enabling the independent latent codes $z_{Y}$ and $z_{S}$ to remain free of unwanted correlations while preserving the predictiveness of the model [6, 20].
Figure 2: Illustration of the data flow. The orange lines connect the information in the observed space and their corresponding latent codes.
From a causal perspective, if $Y$ and $S$ are conditionally independent given $z_{R}$ (i.e., $Y\perp S\mid z_{R}$), $z_{R}$ effectively serves as the common cause, promoting the independence of $z_{Y}$ and $z_{S}$. Figure 2 illustrates this information flow, demonstrating how $z_{R}$ manages correlated information while maintaining the independence of the other latent codes.
Our fairness objective is to achieve independence between $z_{Y}$ and $z_{S}$ by enforcing conditional independence between $\hat{Y}$ and $\hat{S}$ given $z_{R}$, where: $\hat{Y}\!=\!f_{y}(z_{Y},z_{R})$ and $\hat{S}\!=\!f_{s}(z_{S},z_{R})$. Additionally, we utilize $z_{X}$ as an auxiliary latent code to filter out irrelevant features, allowing for more controlled counterfactual generation.
We first consider enforcing the proposed disentanglement by directly imposing Conditional Independence (CI), namely
$p_{\theta}(\hat{Y},\hat{S}\mid z_{R})=p_{\theta}(\hat{Y}\mid z_{R})\,p_{\theta}(\hat{S}\mid z_{R}),$ (7)
which implies $\hat{Y}\,\perp\,\hat{S}\;\big{|}\;z_{R}$. Ideally, one could measure how close $\hat{Y}$ and $\hat{S}$ are to being conditionally independent by minimizing the divergence $D\!\bigl{(}p_{\theta}(\hat{Y},\hat{S}\mid z_{R})\,\big{\|}\,p_{\theta}(\hat{Y}\mid z_{R})\,p_{\theta}(\hat{S}\mid z_{R})\bigr{)}$, but this is generally intractable to compute in practice. Instead, we follow standard information-theoretic principles and propose to minimize the Conditional Mutual Information (CMI): $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$, as a tractable surrogate objective. By definition of CMI,
$I_{\phi}(\hat{Y};\hat{S}\mid z_{R})=$ (8)
$\int\!D_{\mathrm{KL}}\!\Bigl{(}p_{\theta}(\hat{Y},\hat{S}\!\mid\!z_{R})\,\big{\|}\,p_{\theta}(\hat{Y}\!\mid\!z_{R})\,p_{\theta}(\hat{S}\!\mid\!z_{R})\Bigr{)}\,\!dP_{z_{R}},$
and thus minimizing $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ is equivalent to driving the above divergence to zero. See Appendix A for a detailed proof of the equivalence between vanishing conditional mutual information and conditional independence.
From information theory, we know $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})=I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$, thus, when $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ (and equivalently $I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$) is minimized, any undesired dependence between $\hat{Y}$ and $\hat{S}$ is removed.
3.3 Direct Minimization of Conditional Mutual Information
To achieve CI as (7), we propose directly minimizing:
$\min_{\phi}\Bigl{[}I_{\phi}(\hat{Y};\hat{S}\mid z_{R})+I_{\phi}(\hat{S};\hat{Y}\mid z_{R})\Bigr{]},$ (9)
where:
$I_{\phi}(\hat{Y};\hat{S}\mid z_{R})=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid\hat{S},z_{R}),$ (10)
$H_{\phi}(*\mid*)$ stands for the conditional entropy. Incorporating $\hat{S}\!=\!f_{s}(z_{S},z_{R})$, we have:
$I_{\phi}(\hat{Y};\hat{S}\mid z_{R})$ $=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid\hat{S},z_{R})$ (11)
$=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid f_{s}(z_{S},z_{R}),z_{R})$
$=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid z_{S},z_{R})$
$=I_{\phi}(\hat{Y};z_{S}\mid z_{R}),$
from the same transformation (see detailed derivation in Appendix B):
$I_{\phi}(\hat{S};\hat{Y}\mid z_{R})$ $=I_{\phi}(\hat{S};z_{Y}\mid z_{R})$
$=H_{\phi}(\hat{S}\mid z_{R})-H_{\phi}(\hat{S}\mid z_{Y},z_{R}).$ (12)
Therefore, with the introduction of the correlated latent code $z_{R}$ that captures all relevant information between $\hat{Y}$ and $\hat{S}$:
$\min_{\phi}\Bigl{[}$ $I_{\phi}(\hat{Y};\hat{S}\mid z_{R})+I_{\phi}(\hat{S};\hat{Y}\mid z_{R})\bigr{]}$ (13)
$\equiv\min_{\phi}\Bigl{[}I_{\phi}(\hat{Y};z_{S}\mid z_{R})+I_{\phi}(\hat{S};z_{Y}\mid z_{R})\Bigr{]}.$
For the minimization of $I_{\phi}(\hat{Y};z_{S}\mid z_{R})$, as shown in (11) , since $z_{R}$ is given as a condition, we can consider this CMI formula as a function where the independent variable is $z_{S}$ and the dependent variable is $\hat{Y}$, as shown as:
$\mathcal{L}_{\hat{Y}}(z_{S})=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid z_{S},z_{R}),$ (14)
where $z_{R}$ is determined here, $H_{\phi}(\hat{Y}\mid z_{R})$ is a constant, henceforth we need to minimize $-H_{\phi}(\hat{Y}\mid z_{S},z_{R})$. Empirically, we directly minimize the lower bound of it: $-H_{\phi}(\hat{Y}\mid z_{S})$, since: $-H_{\phi}(\hat{Y}\mid z_{S},z_{R})\geq-H_{\phi}(\hat{Y}\mid z_{S})$. Symmetrically, the minimization of $I_{\phi}(\hat{S};z_{Y}\mid z_{R})$ shown in (3.3) is the same concept, see Appendix B for detailed derivation. In this optimization process, $z_{R}$ is responsible for containing any correlation information between target attribute $Y$ and sensitive attribute $S$.
After these simplifications, we introduce the CMI loss to minimize (9):
$\min_{\phi}\Bigl{[}\mathcal{L}_{\text{CMI}}(\omega_{y\_op};\omega_{s\_op};\phi)\Bigr{]}\!\!=\!-(H_{\phi}(\hat{Y}\mid z_{S})+H_{\phi}(\hat{S}\mid z_{Y})),$ (15)
where only update encoder parameters $\phi$. Utilizing opponent classifier $f_{y\_op}(z_{S})$, the entropy term calculation are shown as below:
$H_{\phi}(\hat{Y}\mid z_{S})$ $\!=\!\mathbb{E}_{q_{\phi}(z_{S}\mid x)}\!\biggl{[}-\!\sum_{\,\hat{y}\in\mathcal{Y}}\,\!p_{\theta}(\hat{y}\mid z_{S})\,\!\log\,p_{\theta}(\hat{y}\mid z_{S})\biggr{]}$ (16)
$\!=\!\frac{1}{\lvert B\rvert}\,\!\sum_{i=1}^{\lvert B\rvert}\,\!\sum_{\,\hat{y}\in\mathcal{Y}}\,\!\biggl{[}-\!p_{\theta}\bigl{(}\hat{y}\mid z_{S}^{(i)}\bigr{)}\!\log p_{\theta}\bigl{(}\hat{y}\mid z_{S}^{(i)}\bigr{)}\biggr{]},$
where $p_{\theta}\bigl{(}\hat{y}\mid z_{S}^{(i)}\bigr{)}\!=\!f_{y\_op}\bigl{(}z_{S}^{(i)}\bigr{)}.$ Here, $z_{S}^{(i)}$ denotes the $z_{S}$ sample from the $i$-th element in a mini-batch of size $|B|$; the distribution $q_{\phi}(z_{S}\mid x)$ is given by the encoder. Similar calculation to $H_{\phi}(\hat{S}\mid z_{Y})$, see Appendix C for completed calculation formula. During this optimization process, the opponent classifier parameters $\omega_{y\_op}$ and $\omega_{s\_op}$ are frozen.
3.4 Learning Relevance Between Target And Sensitive Information
To encourage $z_{R}$ capture and only capture the shared information relevant to target property and sensitive property, as well as $z_{Y}$, $z_{S}$ capture main information of $Y$ and $S$ attributes respectively, we propose to maximize the conditional mutual information as Learning Relevance Information loss:
$\min_{\phi}\Bigl{[}\mathcal{L}_{\text{LRI}}(\omega_{y};\omega_{s};\phi)\Bigr{]}\!=\!-(I_{\phi}(\hat{Y};Y\mid z_{R})+I_{\phi}(\hat{S};S\mid z_{R})),$ (17)
where:
$I_{\phi}(\hat{Y};Y\mid z_{R})=H_{\phi}(\hat{Y}\mid z_{R})-H_{\phi}(\hat{Y}\mid Y,z_{R}),$ (18)
$I_{\phi}(\hat{S};S\mid z_{R})=H_{\phi}(\hat{S}\mid z_{R})-H_{\phi}(\hat{S}\mid S,z_{R}),$ (19)
$H_{\phi}(*\mid*)$ stands for the conditional entropy. Maximizing $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{S}\mid z_{R})$ avoid $z_{R}$ capture all information of $Y$ and $S$: $Y\cap S$, which will lead to the Information Bottleneck phenomenon[15, 6, 19] i.e $z_{R}$ capture all the information about target attribute $Y$ and sensitive attribute $S$, degenerating disentanglement performance. On the other hand, minimizing $H_{\phi}(\hat{Y}\mid Y,z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ encourage $z_{R}$ capture relevant information between $Y$ and $S$, which is achieved by marginalizing latent codes in different attribute groups as shown in next.
For entropy calculation of $H_{\phi}(\hat{Y}\mid z_{R})$, We approximate $p_{\theta}(\hat{y}\mid z_{R})$ by marginalizing over $z_{Y}$:
$p_{\theta}(\hat{y}\mid z_{R}^{(k)})$ $=\mathbb{E}_{p(x)}\left[\mathbb{E}_{q_{\phi}(z_{Y}\mid x)}\left[p_{\theta}(\hat{y}\mid z_{Y},z_{R}^{(k)})\right]\right]$ (20)
$\approx\frac{1}{|B|}\sum_{i=1}^{|B|}p_{\theta}(\hat{y}\mid z_{Y}^{(i)},z_{R}^{(k)}),$
where $p_{\theta}(\hat{y}\mid z_{Y}^{(i)},z_{R}^{(k)})\!=\!f_{y}\bigl{(}z_{Y}^{(i)},z_{R}^{(k)}\bigr{)}$, then
$H_{\phi}(\hat{Y}\mid z_{R})\!=\!\mathbb{E}_{q_{\phi}(z_{R}\mid x)}\left[-\sum_{\hat{y}\in\mathcal{Y}}p_{\theta}(\hat{y}\mid z_{R})\log p_{\theta}(\hat{y}\mid z_{R})\right]$ (21)
$\!=\!\frac{1}{|B|}\sum_{i=1}^{|B|}\sum_{\hat{y}\in\mathcal{Y}}\left[-p_{\theta}(\hat{y}\mid z_{R}^{(i)})\log p_{\theta}(\hat{y}\mid z_{R}^{(i)})\right].$
As for the calculation of conditional entropy term $H_{\phi}(\hat{Y}\mid Y,z_{R})$, we regard known condition $Y$ as attribute to grouping samples in a mini-batch of size $|B|$, and calculate the entropy term by marginalizing over $z_{Y}$ within each group, $p_{\theta}(\hat{y}\mid z_{R},y)$ can be computed for $z_{R}^{(k)}$ sampled from an instance $x^{(k)}\in B_{y}$ as:
$p_{\theta}(\hat{y}|z_{R}^{(k)},y)$ $=\mathbb{E}_{p(x\mid Y=y)}\left[\mathbb{E}_{q_{\phi}(z_{Y}\mid x)}\left[p_{\theta}(\hat{y}\mid z_{Y},z_{R}^{(k)})\right]\right]$ (22)
$\approx\frac{1}{|B_{y}|}\sum_{i=1}^{|B_{y}|}p_{\theta}(\hat{y}|z_{Y}^{(i)},z_{R}^{(k)}),$
where $B_{y}$ denotes a subset of the batch with $Y=y$. Then the conditional entropy can be computed as:
$H_{\phi}(\hat{Y}\mid Y,z_{R})=$ (23)
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\!\left[\!\mathbb{E}_{q_{\phi}(z_{R}\mid x)}\!\left[\!-\!\sum_{\hat{y}\in\mathcal{Y}}\!p_{\theta}(\hat{y}\!\mid\!z_{R},y)\!\log p_{\theta}(\hat{y}\!\mid\!z_{R},\!y)\right]\!\right]\!$
$=\frac{1}{|B|}\!\sum_{y\in\mathcal{Y}}\!\sum_{i=1}^{|B_{y}|}\!\sum_{\hat{y}\in\mathcal{Y}}\left[-p_{\theta}(\hat{y}|z_{R}^{(i)},y)\!\log p_{\theta}(\hat{y}|z_{R}^{(i)},y)\right].$
The calculation of $H_{\phi}(\hat{S}\mid z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ are similar to $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{Y}\mid Y,z_{R})$ respectively, see Appendix C for completed calculation formula. Plugging these estimates(21)(LABEL:cal_H_Y_Y_zr) back into (17), shared feature between $Y$ and $S$ will be learned in $z_{R}$ while getting rid of Information Bottleneck phenomenon. Note that the classifier parameters $\omega_{y}$ and $\omega_{s}$ are remain freeze when (17) optimizing.
3.5 Final Objective Function
To integrate the above components into a coherent training framework, we employ the two-step optimization strategy defined in (24) and (25).
$\min_{\theta,\phi,\omega_{y},\omega_{s}}\Bigl{[}\mathcal{L}_{\text{VAE}}(\theta,\phi)+\bigl{(}\mathcal{L}_{y}(\omega_{y},\phi)+\mathcal{L}_{s}(\omega_{s},\phi)\bigr{)}\Bigr{]}\!\!+\!\min_{\phi}\Bigl{[}$ (24)
$\lambda_{\text{CMI}}\mathcal{L}_{\text{CMI}}(\omega_{y\_op},\omega_{s\_op};\phi)\!+\!\mathcal{L}_{\text{TC}}(\phi)\!+\!\lambda_{\text{LRI}}\,\mathcal{L}_{\text{LRI}}(\omega_{y},\omega_{s};\phi)\Bigr{]}$
Specifically, in (24), we jointly update $(\theta,\phi,\omega_{y},\omega_{s})$ by minimizing the VAE loss (2) alongside the main classification losses (3) and (4), which together reformulate the ELBO. We further include the CMI loss (15) to reduce unwanted information leakage, the LRI loss (17) to capture shared patterns in $z_{R}$, and the TC penalty (1) to promote factorization among the latent codes $(z_{Y},z_{R},z_{S})$.
$\min_{\omega_{y\_op},\omega_{s\_op}}\Bigl{[}\mathcal{L}_{y\_op}(\omega_{y\_op};\phi)+\mathcal{L}_{s\_op}(\omega_{s\_op};\phi)\Bigr{]}$ (25)
In parallel, the second procedure (25) optimizes $(\omega_{y\_op},\omega_{s\_op})$ by minimizing the opponent classification losses (5) and (6) while holding $\phi$ fixed.
The hyperparameters $\lambda_{CMI},\lambda_{LRI}>0$ control the relative importance of these terms, ensuring each network component learns its designated function while enforcing minimal information leakage, preserving shared information in $z_{R}$ and maintaining the salient factors for $Y$ and $S$ in $z_{Y}$ and $z_{S}$ respectively. See hypermeter analysis in Appendix D.
4 Experiment
Figure 3: Examples of Fair Counterfactual Generation. The first row shows the source and reference images. Rows 2–5 display counterfactuals obtained by replacing latent subspaces $z_{X}$, $z_{Y}$, $z_{S}$, and $[z_{S},z_{R}]$, respectively. Notably, the replacement with $[z_{S},z_{R}]$ (row 5) naturally adapts sensitive features for different sensitive attributes without domain knowledge. (mustache for men and makeup for women).
To ensure a rigorous and comprehensive evaluation, we conduct experiments comparing our proposed method with a diverse set of state-of-the-art approaches across multiple categories of learning paradigms in various tasks. Specifically, we include FairFactorVAE [28], FairDisCo [27], FFVAE [6], GVAE [7], ODVAE [42] and FADES [15], shown in Section 2.1. As for traditional correlation-aware learning that is discussed in Section 2.2, since it require additional annotated data to build causal graph, we except them in our experiment. Further details on the experimental setup and additional discussions are provided in Appendix D.
4.1 Fair Classification
Table 1: Evaluation of downstream classification tasks on various datasets from learned representation. Best in bold, second in red.
                 Downstream Classification Performance Methods CelebA [29] UTKFace [49] Dogs and Cats [37] Color bias MNIST [17] Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$   FADES [15] [CVPR’24] 0.918 0.034 0.135 0.812 0.059 0.139 0.769 0.058 0.086 0.973 0.094 0.160 GVAE [7] [CVPR’20] 0.919 0.047 0.131 0.819 0.204 0.197 0.748 0.064 0.131 0.961 0.109 0.176 FFVAE [6] [PMLR’19] 0.892 0.076 0.072 0.766 0.269 0.201 0.729 0.059 0.110 0.952 0.081 0.092 ODVAE [42] [ECCV’20] 0.886 0.039 0.103 0.736 0.165 0.210 0.689 0.051 0.038 0.957 0.247 0.162 FairDisCo [27] [KDD’22] 0.839 0.074 0.051 0.766 0.266 0.200 0.680 0.115 0.111 0.949 0.129 0.136 FairFactorVAE [28] 0.914 0.055 0.136 0.720 0.096 0.134 0.707 0.055 0.110 0.957 0.096 0.128 CAD-VAE (Ours) 0.923 0.021 0.112 0.817 0.045 0.137 0.773 0.048 0.090 0.977 0.076 0.118
The objective of fair classification is to achieve a balance between minimizing fairness violations and maintaining high predictive performance. To evaluate the effectiveness of our proposed method, we conduct experiments on a diverse set of benchmark fairness datasets. For facial attribute classification tasks, we utilize the CelebA [29] and UTKFace [49] datasets. Following prior works [44, 47, 48, 15], we set the CelebA classification task to predict the “Smiling” attribute, while for UTKFace, the objective is to classify whether a person depicted in the image is over 35 years old, with gender serving as the sensitive attribute. Additionally, the Dogs and Cats dataset [37] is used to distinguish between dogs and cats, with fur color as the sensitive attribute. Furthermore, we assess fair classification performance using the Colored MNIST dataset [18, 17, 35], which incorporates a controlled color bias in the standard MNIST dataset to simulate spurious correlations. To assess fairness violations, we use standard metrics including Demographic Parity (DP) [2] and Equalized Odds (EOD) [10]. See detailed experimental setup in Appendix D.
The result of fair classification can be seen in Table 1. Across all evaluated datasets, our method consistently achieves state-of-the-art classification accuracy and fairness, validating its effectiveness in robust disentanglement by preserving high-quality target-related information while minimizing sensitive attribute leakage. Among the baselines, FADES achieves competitive accuracy and strong fairness. However, its inability to strictly minimize conditional mutual information results in sensitive information leakage and an excessive capture of target-related data in the relevance latent code $z_{R}$, which limits classification performance. In contrast, our approach directly minimizes the conditional mutual information between $z_{Y}$ and $z_{S}$ regarding their respective opposite attributes $S$ and $Y$, and incorporates an explicit relevance learning strategy, as detailed in Section 3.3 and Section 3.4. This design ensures that $z_{R}$ captures sufficient, but not excessive, information about $Y$ and $S$, thereby enhancing fairness while preserving essential target information. More fair classification experiment results can be seen in Appendix E.
4.2 Fair Counterfactual Generation
We evaluate our approach on the CelebA dataset [29], a widely-used benchmark for facial attribute manipulation. We select Smiling as the target label $Y$ and Gender as the sensitive attribute $S$. In our experiments, we substitute specific latent code of source images with reference images, including $z_{X}$, $z_{Y}$, $z_{S}$, and $[z_{S},z_{R}]$. Figure 3 illustrates the generated counterfactuals, with the first row showing source and reference images and subsequent rows demonstrating the effects of substituting each latent code.
The experimental results demonstrate the effectiveness of our method in generating fair counterfactuals. As shown in Figure 3, substituting $[z_{S},z_{R}]$ (Row 5) leads to a natural adaptation of sensitive-relevant features without domain-specific knowledge. For instance, the model automatically adds makeup to female images and a mustache to male images, highlighting the semantic alignment of $z_{R}$ with both the target and sensitive attributes. Compared to substituting only $z_{S}$, our approach achieves more interpretable translation, ensuring that fairness is maintained throughout the counterfactual generation. More fair counterfactual generation experiment results can be seen in Appendix E.
To quantitatively assess the quality of the generated counterfactuals, we compare evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained by randomly permuting $z_{Y}$ and $z_{S}$ within the evaluation set. Specifically, we use the FID [12, 15] to assess reconstruction fidelity and the Inception Score (IS) [5] to evaluate semantic and perceptual quality. Lower $\Delta\text{FID}$ values indicate minimal distortion and higher translation quality, while lower $\Delta\text{IS}$ values suggest that semantic and perceptual attributes are well preserved. Detailed experimental settings are provided in Appendix D.
Table 2: FID and IS difference between original reconstruction and perturbed target/sensitive codes’ reconstruction.
  CAD-VAE FADES GVAE FFVAE ODVAE FairFactorVAE $\Delta\textit{\text{FID}}\downarrow$ 1.072 1.167 3.710 1.409 14.647 6.239 $\Delta\textit{\text{IS}}\downarrow$ 1.214 2.379 3.148 3.829 6.113 5.378
Quantitative analysis in Table 2 further validates our approach. Our method achieves both lower $\Delta\text{FID}$ and $\Delta\text{IS}$ compared to other fair representation learning methods, demonstrating that our fair counterfactual generation approach renders counterfactuals with superior image quality and minimal distortion.
4.3 Fair Fine-Grained Image Editing
With the introduction of the correlated latent code $z_{R}$, fair fine-grained image editing—as a fundamental concept in counterfactual fairness—can be naturally achieved by aligning latent codes from different samples. We use linear interpolation to synthesize a latent code: $z^{\prime}=(1-\lambda)z_{1}+\lambda z_{2}$, where $z_{1}$ is the latent code from the source image and $z_{2}$ is the corresponding code from the reference image. The synthesized $z^{\prime}$ replaces $z_{1}$, enabling a gradual transfer from the source to the reference latent code.
In our experiments, following the setup in Section 4.2 where Smiling is the target label $Y$ and Gender is the sensitive attribute $S$, we generate interpolated latent codes between source and reference images. In the blue-framed subfigure of Figure 1, images are generated by interpolating $z_{Y}$ and $z_{S}$: the horizontal axis shows the transition of $z_{S}$ from the source to the reference image, while the vertical axis shows the corresponding change in $z_{Y}$. During this interpolation, $z_{R}$ and $z_{X}$ remain unchanged, with the interpolation parameters for both $z_{Y}$ and $z_{S}$ set to $\lambda\in\{0,0.33,0.66,1\}$.
Similarly, in the red-framed subfigure, images are generated by interpolating $z_{Y}$ and $z_{R}$. Here, the horizontal axis corresponds to the transition of $z_{R}$ from the source to the reference image, and the vertical axis corresponds to $z_{Y}$. Note that the source image’s $z_{S}$ is fully replaced by that of the reference image, and $z_{X}$ remains constant. Initially, the interpolation parameters for $z_{R}$ are set to $\lambda\in\{0.5,1\}$, and when combined with the final column of the blue-framed subfigure, the range is extended to $\lambda\in\{0,0.5,1\}$.
Figure 1 demonstrates a smooth transformation of each attribute, with modifications in one latent code minimally affecting the others, a key characteristic of effective disentanglement. Specifically, as the correlated latent code $z_{R}$ captures sensitive relevant information, we can explicitly control these properties: in the left subfigure, we gradually introduce makeup (such as enhanced lipstick and eyeshadow), while in the right subfigure, we progressively add a mustache. More fair fine-grained image editing experiment results can be seen in Appendix E.
Similarly, we measure $\Delta\text{FID}$ and $\Delta\text{IS}$ to quantitatively assess the quality of fine-grained image editing. Unlike the evaluation setup in Section 4.2, we compute the differences in evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained through latent code traversals for each $\lambda$ combination. Detailed experimental settings are provided in Appendix D. Table 3 summarizes the comparison results, showing that our method exhibits both lower $\Delta\text{FID}$ and $\Delta\text{IS}$ values compared to other fair generation methods. These results indicate that our fine-grained image editing approach not only ensures smoother attribute transformations and superior image fidelity, but also allows for more precise control of task-relevant features.
4.4 Fair Text-to-Image Editing
(a) Original
(b) StyleCLIP
(c) StyleCLIP+
Figure 4: Style transfer using StyleCLIP and the CAD-VAE extension. This example transforms the (a) into “a dancer with long blonde hair.” “StyleCLIP+” means StyleCLIP + CAD-VAE.
To further validate the capability and explore the applicability of our method, we integrated it as an adaptor on top of a pre-trained, frozen CLIP image encoder [39] and trained it on Facet dataset [9] to enhance fairness in vision-language tasks. Table 4 presents the experimental results. These results demonstrate that our approach significantly improves fairness without compromising performance compared to the linear probing baseline (ERM), underscoring its potential for a range of vision-language tasks with fairness considerations.
Furthermore, we applied our method in StyleCLIP [38] as a fair discriminator to address inherent fairness issues, such as career-gender biases, which persist even when an identity preservation loss is employed. As illustrated in Figure 8, StyleCLIP [38] exhibits a bias by correlating the role of “dancer” with a specific gender. In contrast, our method effectively mitigates this bias while maintaining the efficacy of attribute modification. See Appendix E for details.
Table 3: FID and IS difference between original reconstruction and traversed target/sensitive codes’ reconstruction.
  CAD-VAE FADES GVAE FFVAE ODVAE FairFactorVAE $\Delta\textit{\text{FID}}\downarrow$ 1.642 2.362 4.023 2.789 15.893 7.120 $\Delta\textit{\text{IS}}\downarrow$ 1.849 2.919 4.848 5.292 6.890 5.767
Table 4: Performance of CLIP(ViTB/32) on Facet dataset. WG: Worst Group, Gap: Difference between WG and Avg.
  Method Top-1 Acc. (%) Top-3 Acc. (%) WG $\uparrow$ Avg $\uparrow$ Gap $\downarrow$ WG $\uparrow$ Avg $\uparrow$ Gap $\downarrow$   Zero-shot 2.79 53.45 50.66 15.31 76.79 61.48 Linear prob 1.17 65.46 64.29 1.79 85.34 83.55 CAD-VAE 69.97 70.54 0.57 85.36 85.95 0.59
5 Conclusion
Our method aims to solve fairness concerns in representation learning and deep generative models. By introducing a correlated latent code that captures shared information, sensitive information leakage can be eliminated directly and efficiently without conflicting with the prediction objective, which is a core issue in disentanglement, by minimizing the conditional mutual information between target latent code and sensitive latent code. Parallel with our explicit relevance learning strategy imposed on the correlated latent code, it is encouraged to capture the essential shared information that cannot be perfectly separated without additional domain knowledge. Various benchmark tasks further demonstrate the robustness and wide applicability of our method. Further details on the experimental setup, results, discussion, and potential future directions are provided in the Appendix.